OpenStudy (sloppycanada):

Determine the linear velocity of an object with an angular velocity of 3.7 radians per second at a distance of 10 centimeters from the center.

OpenStudy (kanwal32):

3.7*.1

OpenStudy (sloppycanada):

So just 37? There's no equation? There's usually the following equation involved - v = rw

OpenStudy (kanwal32):

yes

OpenStudy (kanwal32):

convert cm to m

OpenStudy (kanwal32):

for m/s

OpenStudy (sloppycanada):

Huh... then why give the equation?

OpenStudy (anonymous):

r is 0.1m and w is 3.7rad/s. plug it in the equation

OpenStudy (sloppycanada):

And that's 1.162...

OpenStudy (sloppycanada):

@Astrophysics If you get a sec.

OpenStudy (astrophysics):

\[v = \omega r \] omega is your angular velocity, v is your linear velocity, and r is your radius

OpenStudy (sloppycanada):

So it's v = 10(3.7pi)

OpenStudy (sloppycanada):

But that's not one of my available choices.

OpenStudy (astrophysics):

Because you have to pay attention to the units

OpenStudy (sloppycanada):

So... 10 centimeters?

OpenStudy (sloppycanada):

Do I divide something?

OpenStudy (astrophysics):

What do you want your answer in

OpenStudy (sloppycanada):

I'm sorry, I've never been good with units.

OpenStudy (sloppycanada):

Centimeters per second?

OpenStudy (sloppycanada):

If I want seconds, do I convert the radians to seconds?

OpenStudy (astrophysics):

Yes then we just multiply to get the following \[v = 10 cm \times 3.7 \frac{ rad }{ s } = 37 cm/s\]

OpenStudy (sloppycanada):

How would I do that in a calculator though?

OpenStudy (astrophysics):

?

OpenStudy (sloppycanada):

rad/s = pi/60?

OpenStudy (astrophysics):

\[1 ~ radian = \frac{ 180 }{ \pi } degrees\] these are separate units we don't have to do anything we just have to multiply the angular speed by the radius

OpenStudy (sloppycanada):

Oh so I have to convert it to degrees?

OpenStudy (astrophysics):

No, you don't need to convert anything

OpenStudy (sloppycanada):

Then I am absolutely confused.

OpenStudy (astrophysics):

The radians is an "invisible" unit

OpenStudy (astrophysics):

So if we have rad/s*m = m/s

OpenStudy (sloppycanada):

So for this one... Determine the linear velocity of an object with an angular velocity of 5.9 radians per second at a distance of 12 centimeters from the center. 5.9 x 60?

OpenStudy (astrophysics):

Where did you get 60?

OpenStudy (sloppycanada):

60 seconds?

OpenStudy (sloppycanada):

If I follow the whole v = rw I get some insane answer that has nothing to do with my choices.

OpenStudy (astrophysics):

\[\huge v = \omega r \] your linear velocity here is v, omega (w) is the angular velocity and r is the radius

OpenStudy (astrophysics):

Why are you using 60 seconds, it's just \[v = (5.9 rad/s) (12 cm)\]

OpenStudy (astrophysics):

That gives your answer in cm/s

OpenStudy (sloppycanada):

But that's like 222.43

OpenStudy (astrophysics):

Use your calculator, 5.9*12 = 70.8

OpenStudy (sloppycanada):

Oh.. so I just get rid of the pi

OpenStudy (astrophysics):

Why are you using pi??

OpenStudy (sloppycanada):

I don't know! I meant the radians symbol thing

OpenStudy (astrophysics):

 all we need is that relationship